In the table, you find figures taken from some Country’s national accounts for the 2018 fiscal year. Figures are in billions of dollars.

Find:

a) Savings

b) Private Savings

c) Net Government Income

d) Gross National Product

e) Net Exports

f) Government Interest Payments

g) Current Account Balance

Variables: Values:
Y(GDP) 2000
NFP (Net Factor Payment) 100
C (Consumption) 1400
G (Government Expenditure) 200
PDI (Personal Disposable Income) 1600
I (Investment) 160
T (Taxes) 640
TR (Government Transfers) 130
Answer #1

A. Saving = Disposable Income - Consumption

Saving = 1600 - 1400

Saving = $ 200 billion

B. Private saving = National Income - Consumption - Taxes

Private saving = 2,000 - 1,400 - 640

Private Saving = - $ 40 billion

C. Net government income =Consumption +investment +government spendimg+ net export + Net factor income - indirect tax - depreciation

Net government income = 1400 + 160 + 200 + 100 -640

Net government income = 1220

D. GNP = GDP + Net factor payment = 2100

E. Net export = GDP - consumption - government spending- investment

Net exports = 2000 - 1400 - 200 - 160

Net export = 240

F. Government interest payment = 40

G. Current account balance = net export + at income from abroad + current transfer

Current account balance = 240 + 100 + 130 = 370